OpenStudy (anonymous):

Let h(x) = 2x + 7. Find the value of h(x) + 14.

OpenStudy (hoblos):

2x + 21

OpenStudy (anonymous):

How did you get to that answer?

OpenStudy (mr.math):

You have \(h(x)=2x+7 \implies h(x)+14=(2x+7)+14=2x+(7+14)=2x+21.\)

OpenStudy (hoblos):

2x + 7 + 14 =2x + 21